OpenStudy (anonymous):

1and1over4 times 3over 5 is show work or not

OpenStudy (anonymous):

use equation editor or type using parenthesis

OpenStudy (anonymous):

\[1\tfrac{1}{4}\times \frac{3}{5}=\frac{5}{4}\times \frac{3}{5}=\frac{3}{4}\]